Calculate the pH of the solution that results from
Answer #1

a)

this is abuffer so

pH = pKa + log(NaCHO2/HCHO2)

NaCHO2 = MV = 55*0.15 = 8.25

HCHO2 = MV = 70*0.13 = 9.1

pH = 3.75 + log(8.25/9.1) = 3.707

B)

pOH = pKB + log(NH4/NH3)

NH4= MV = 0.13*120 = 15.6

NH3 = MV = 0.13*260 = 33.8

pOH = 4.75 + log(15.6/33.8 )=5.9431

pH = 14-5.9431

pH = 8.0569